OpenStudy (anonymous):

AN airplane flies 530km against the wind and 670 km with the wind in a total time of 4hr. The speed of the airplane in still air is 300km/h. What is the speed of the wind?

OpenStudy (anonymous):

you first need to find the km/h for 670km in 4 hours

OpenStudy (anonymous):

r*t=d Let s and w be the speed of the plane in still air and w be the speed of the wind. Sole the following equations for w:\[\{4(s-w)=530,4(s+w)=670\}\]\[w=\frac{35}{2} km/hr\]